Two bodies x and y are projected on the same horizontal plane with same initial speed. Unit angle 30° and 60° respectively to the horizontal. Neglecting air resistance, the ratio of the range of the range of x to that of y is?



1
Expert's answer
2020-12-08T10:37:11-0500
"R_x=\frac{v^2\sin{2(30)}}{g}=\frac{v^2\sin{60}}{g}\\
R_y=\frac{v^2\sin{2(60)}}{g}=\frac{v^2\sin{120}}{g}\\
\frac{R_x}{R_y}=\frac{\sin{60}}{\sin{120}}=1"